∫(1- x 2 /x-2x 2 )dx

∫(1-x/x-2x2)dx

Grade:12

1 Answers

Vikas TU
7 years ago
Do it by partial fraction,
∫(1-x/x-2x2)dx can be written as:
(1-x)*(1 +x)/x*(1 -2x)dx
= A/x + B/(1 -2x)
(1 -x) (1 +x) =A(1- 2x) + B(x)
Put x= 1, -1 and get the values of A nd B and then integerate.
Answer will ome in lagarith form.